3x+14y=2 and 6x+8y=4 solve by elimination
Answer by checkley77(7072)

3x+14y=2 multiply by -2 & add.
6x+8y=4
-6x-28y=-4
-------------
-20y=0
y=0 ans.
3x+14*0=2
3x=2
x=2/3 ans.
6*2/3+8*0=4
12/3=4
4=4